# GPT vs RETRO: Exploring the Intersection of Retrieval and Parameter-Efficient Fine-Tuning

### v3

@@ 4.1 Main Results
3) P-tuning underperforms LoRA and Adapters in smaller GPT models but bests them in larger sizes. This difference is visualized in Figure 3 and Figure 3 (Appendix B). However, for RETRO models, P-tuning generally under performs the other PEFT methods across all model sizes. We believe that P-Tuning’s lower parameter count contributes to its lower performance especially when paired with smaller base model sizes. For RETRO specifically P-Tuning We hypothesis that P-Tuning’s weaker ability in all RETRO model sizes could lie in architecture differences. In P-tuning, virtual tokens are intentionally prepended to the decoder’s input, but they are not included in the retrieval encoder. Although they can influence the encoder through cross-attention, the impact might not be as direct or substantial as required. Alternatively, LoRA and Adapters are added to both encoder and decoder which explains their improved capabilities.
4) The performance ceiling for PEFT-tuned models is notably higher for GPT than RETRO. This is demonstrated in Figure 4 (Appendix B) where example, using medium-sized models, the average score of LoRA with GPT is 40.85, while with RETRO it is 34.24. This disparity suggests that GPT has more room for improvement with PEFT tuning. This phenomenon can also be possibly explained by the two different pre-training strategies. Since GPT pre-training is not focused on retrieval-augmented generation, it opens larger room for improvement during fine-tuning.

### v4

@@ 4.1 Main Results
3) P-tuning underperforms LoRA and Adapters in GPT and RETRO models for these tasks. This difference is visualized in Figure 3 and Figure 4 (Appendix B). For RETRO models, P-tuning under performs the other PEFT methods more significantly. We believe that P-Tuning’s lower parameter count contributes to its lower performance especially when paired with smaller base model sizes. For RETRO P-Tuning specifically, we hypothesis that P-Tuning’s weaker ability in all RETRO model sizes could lie in architecture differences. In P-tuning, virtual tokens are intentionally prepended to the decoder’s input, but they are not included in the retrieval encoder. Although they can influence the encoder through cross-attention, the impact might not be as direct or substantial as required. Alternatively, LoRA and Adapters are added to both encoder and decoder which explains their improved capabilities.
4) The performance ceiling for PEFT-tuned models is notably higher for GPT than RETRO. This is demonstrated in Figure 5 (Appendix B) where for example, with medium-sized models, the average score of LoRA with GPT is 40.85, while with RETRO it is 34.24. This disparity is seen in all the PEFT methods across multiple model sizes and suggests that GPT is better suited for PEFT. This phenomenon can also be possibly explained by the two different pre-training strategies. Since GPT pre-training is not focused on retrieval-augmented generation, it allows for a larger room for improvement when fine-tuned on these tasks.